### v2

# Feature Calibration enhanced Parameter Synthesis for CLIP-based Class-incremental Learning

Abstract
Class-incremental Learning (CIL) enables models to continuously learn new class knowledge while memorizing previous classes, facilitating their adaptation and evolution in dynamic environments. Traditional CIL methods are mainly based on visual features, which limits their ability to handle complex scenarios. In contrast, Vision-Language Models (VLMs) show promising potential to promote CIL by integrating pretrained knowledge with textual features. However, previous methods make it difficult to overcome catastrophic forgetting while preserving the generalization capabilities of VLMs. To tackle these challenges, we propose Feature Calibration enhanced Parameter Synthesis (FCPS) in this paper. Specifically, our FCPS employs a specific parameter adjustment mechanism to iteratively refine the proportion of original visual features participating in the final class determination, ensuring the model’s foundational generalization capabilities. Meanwhile, parameter integration across different tasks achieves a balance between learning new class knowledge and retaining old knowledge. Experimental results on popular benchmarks (e.g., CIFAR100 and ImageNet100) validate the superiority of the proposed method.
Index Terms:
Class-Incremental Learning, Visual Language Learning, Knowledge Distillation, Feature Calibration.
I Introduction
Class-incremental Learning (CIL) aims to incorporate the knowledge of new classes incrementally and build a universal classification model among all seen classes. In other words, CIL should not only continuously acquire knowledge from the current task but also preserve the knowledge from former tasks, making it crucial for real-world applications that require adaptability. For example, autonomous driving systems must dynamically adapt to evolving road conditions, weather variations, traffic regulations, and driving patterns[1], while medical imaging systems face challenges from emerging diseases and evolving pathological presentations. These scenarios require CIL-enabled models that effectively integrate new information without revisiting previous data - a necessity for deployment on edge devices constrained by limited storage and computational capacity.
A primary challenge in CIL is to mitigate catastrophic forgetting[2], where the integration of new knowledge could overwrite previously acquired information in the process of continuously building a classifier for all classes. Effective solutions must balance stability—the retention of past knowledge—and plasticity—the adaptability to new tasks [3]. To address this, various approaches have been proposed, including replay-based[4], regularization-based [5], and parameter-isolated methods[6]. For instance, [4] proposes an efficient memory scheme based on the Resnet-18 to cope with the highly imbalanced data problem, while DNE [7] introduces a dynamic architecture that balances accuracy and model complexity based on Vision Transformer (ViT)[8] backbone. While these strategies have shown success with foundational vision backbones (as shown in Fig. 1(a)) like ResNet[9] and ViT[8], their reliance on visual features alone limits their applicability to complex scenarios.
Figure 1: Illustration of existing CIL framework.
Recent advancements in incremental learning have explored Vision-Language Models (VLMs), such as ViT-BERT [10] and CLIP [11], which harness both visual and textual information to enhance performance. These pretrained models have learned rich feature representation from large amounts of training data[12, 13], which can be used to bridge the knowledge gap between class information trained on different stages. By freezing CLIP’s pretrained backbone and leveraging its zero-shot capabilities, Continual-CLIP[14] demonstrates the pretained VLMs are well suited for CIL. Motivated by the zero-shot ability of CLIP, the follow-up methods propose different finetuning methods to mitigate catastrophic forgetting, achieving superior performance to traditional methods[15].
Prompt-tuning[16, 17] and visual feature tuning [18] are two representative categories of CLIP-based approaches for incremental learning. Prompt-tuning employs task-specific or input-conditioned prompts to guide the model while preserving the original backbone, while visual feature tuning modifies the pretrained model to accommodate new tasks as illustrated in Fig. 1(c). Even though the previous methods can inject valuable pretrained knowledge, they still face challenges for CIL. First, prompt-tuning methods primarily focus on improving and enhancing the text prompt space in Fig. 1(b), overlooking the importance of effectively utilizing visual features. Second, visual feature tuning methods risk degrading its feature extraction ability based on its coverage of the original CLIP features. As shown in Fig. 2, the last-stage classification clusters of visual feature tuning method is not even seperated clearly compared to those of zero-shot CLIP. The recently proposed RAPF[19] is a visual feature tuning method that uses neighboring categories separation to enhance the representation of the feature space. However, the new and old knowledge is projected into one shared feature basis, limiting the ability to adapt to the shared feature space dynamically and adopting the CLIP features.
Figure 2: T-SNE visualization of (a) Continual-CLIP, (b) visual feature tuning and (c) Our method with CIFAR100 B0 Inc10 after the second task.
To address these issues, this paper proposes a novel CIL method called Feature Calibration enhanced Parameter Synthesis (FCPS). We adopt the frozen encoder of CLIP as the foundational architecture for our approach. Considering the potential distortion of the pretrained knowledge of CLIP caused by fine-tuning, we designed Feature Calibration (FC) which uses the trainable layer to transform the CLIP features into new feature spaces. FC effectively utilizes and rationalizes the enhancement of the original visual features of CLIP while introducing specialized parameters to control the retention of the original visual features of CLIP precisely. It allows the model to maintain strong generalization capabilities while reducing the likelihood of overfitting. At the same time, we focus on whether the model can be dynamically adapted between old and new knowledge in the process of continuously building a classifier for all classes, rather than on the adaptation and improvement of textual cues. We introduce the Parameter Synthesis (PS) module, which effectively mitigates catastrophic forgetting by addressing the significant discrepancies in knowledge between adjacent tasks. PS integrates the adapter parameters of the current stage with those of the previous stage to initialize the adapter parameters for the next stage. With task evolution, the parameters of adjacent stages undergo decomposition and integration, alleviating the forgetting issue as tasks progress. Furthermore, we incorporate an additional dynamic distillation loss during the training process to constrain the changes of adapter parameters, further mitigating forgetting. The main contributions of this paper are as follows:
  • We propose a tuning-efficient VLM-based framework for CIL, which effectively solves the issue of the original feature of CLIP being distorted due to fine-tuning with a feature calibration module.
  • To tackle the catastrophic forgetting problem, we introduce the QR decomposition-based parameter synthesis, which captures parameter changes during task transitions and fuses the parameters constructed for different tasks.
  • Experimental results on CIFAR100 and ImageNet100 datasets show that our method achieves comparable performance to state-of-the-art methods.
Figure 3: The framework of the proposed feature calibration enhanced parameter synthesis.
II Related work
II-A Class-Incremental Learning
CIL aims to continuously incorporate new knowledge from new data while retaining, integrating and even optimizing old knowledge to prevent catastrophic forgetting [3] in the process of continuously building a classifier for all classes. The current mainstream methods for CIL are divided into three main categories. The replay-based methods either save a portion of samples from previous tasks [20] or generate synthetic samples with a generative model [21] in a cache. When training on new tasks, these samples are used alongside new samples to prevent forgetting previously learned knowledge. Regularization-based methods impose constraints on network parameter updates. Weight regularization focuses on preventing drift in weights associated with previous tasks [5], while data regularization methods transfer knowledge from the old model to the new one through knowledge distillation [22]. Parameter isolation-based methods mitigate catastrophic forgetting by dynamically increasing the capacity of network. It enables the model to retain knowledge from previous tasks while expanding to learn new ones [6]. Due to the excellent zero-shot capability of CLIP on single or multi-task [23, 13], recent works [18, 16, 17] have shown the possibility of applying CLIP to incremental learning scenarios.
II-B Vision-Language Models
Large-scale visual language models like CLIP [11] and ALIGN [24] have been pretrained on a large number of images and texts to achieve a unified embedding space across modalities. Due to their high transferability, they can be applied to downstream tasks in a zero-shot fashion. A recent emerging trend in the field of continuous learning is to combine large-scale vision models with additional learnable parameters or modules to adapt the model to downstream tasks. For example, L2P [25] introduces prompt-based learning into continual learning and proposes enabling a single pretrained model to adapt to sequential tasks through a shared pool of prompts. DualPrompt [15] proposes attaching complementary prompts to the pretrained backbone and formulates the objective as learning task-invariant and task-specific instructions. Most of these works focus on prompt adjustment, prefix adjustment and effective organization of text features but ignore the effective attention to visual features. This paper aims to fill this research gap by proposing a comprehensive solution that focuses on visual features to tune visual language models without suffering from forgetting.
III Methodology
III-A Overview
Traditional visual methods cannot effectively utilize textual information and the fine-tuning methods based on CLIP suffer from catastrophic forgetting. We propose a method called FCPS that effectively leverages cross-modal knowledge from CLIP while simultaneously addressing the forgetting issue. As shown in Fig. 3, this method enhances performance from three key perspectives, balancing both retention and comprehensiveness. To reduce the risk of distorting the pretrained CLIP knowledge caused by direct fine-tuning, the Feature Calibration Sec. III-B is designed to optimize and enhance visual features. After this process, a special parameter $\alpha$ is introduced to retain a portion of the image embedding features obtained from the Image Encoder. It allows the model to maintain good generalization while mitigating the risk of overfitting. Meanwhile, PS effectively integrates the current adapter parameters with those from the previous stage, serving as the initial parameters for the next stage’s adapter. PS continuously updates and iterates as tasks evolve, enabling the effective retention of old class knowledge and alleviating the issue of catastrophic forgetting in the process of continuously building a classifier for all classes. Finally, the introduction of dynamic distillation allows for the effective transfer and integration of old and new knowledge from current and previous tasks, further ensuring that forgetting is mitigated.
III-B Feature Calibration
Given the powerful zero-shot capability of CLIP and broad transferability to downstream tasks, we designed a Feature Calibration mechanism. It specifically integrates task-specific features with the original features extracted by CLIP, ensuring the overall generalization ability of the model. As illustrated in Fig. 3, the image features $\mathbf{F}_{\text{fine-tuned}}$ are obtained by passing the images through the image encoder and adapter, while the label features $\textbf{F}_{\text{text}}$ are generated by feeding class labels into the prompt generator and subsequently into the text encoder. The classification results are determined by measuring the similarity between these features.
We designed a trainable adaptive layer called Enhanced Feature Synthesis (EFS), which consists of four layers of linear transformations and incorporates attention mechanism techniques to achieve feature enhancement. The feature dimensions remain the same before and after transformation by EFS. To retain the original characteristics of CLIP, feature calibration is conducted as follows:
$\textbf{F}_{\text{fine-tuned}}=(1-\alpha)\Phi_{EFS}\left(\textbf{F}_{\text{image }}\right)+\alpha\textbf{F}_{\text{image}},$ (1)
where $\alpha$ is a parameter that balances the enhanced features with the original CLIP knowledge. The first term represents the adaptive learning of knowledge from new categories in the new task, while the second term is specifically designed to retain the original knowledge from CLIP.
III-C Parameter Synthesis
We introduced Parameter Synthesis to integrate the adapter parameters from the current stage with those from the previous stage, enabling effective fusion of new and old knowledge in the process of continuously building a classifier for all classes. As tasks dynamically evolve, this approach effectively mitigates the problem of forgetting.
In incremental learning, traditional weight update methods often face the problem of catastrophic forgetting, especially when the model learns new tasks and the knowledge from old tasks may be forgotten. To address this issue, we propose a method based on QR decomposition, which effectively integrates new and old knowledge by decomposing and recombining the weight matrix, thus reducing the forgetting of old knowledge.
Specifically, QR decomposition decomposes a target matrix into two matrices $\mathbf{Q}$ and $\mathbf{R}$, where $\mathbf{Q}$ is an orthogonal matrix and $\mathbf{R}$ is an upper triangular matrix. This decomposition method effectively preserves the structure of the matrix while ensuring that learning new tasks does not excessively interfere with the knowledge of old tasks. We apply QR decomposition to the old and new weight matrices $\{\mathbf{W}_{t-1},\mathbf{W}_{t}\}$:
$\mathbf{W}_{t-1}=\mathbf{Q}_{t-1}\mathbf{R}_{t-1},\mathbf{W}_{t}=\mathbf{Q}_{t}\mathbf{R}_{t}.$ (2)
We then use a weight coefficient $\beta$ to balance the influence of new and old tasks and fuse the components of the two matrices accordingly. The final fused weight matrix, which becomes the initial weight matrix for the model’s adapter in the next stage, is updated as:
$\mathbf{W}=\mathbf{Q}_{t-1}\mathbf{R}_{t+1},$ (3)
$\mathbf{R}_{t+1}=\left[\mathbf{Q}_{t-1}^{T}\otimes\mathbf{Q}_{t}\cdot\beta\right]\oplus\left[\mathbf{R}_{t-1}\cdot(1-\beta)\right],$ (4)
where $\beta$ is adjusted to maximize the retention of knowledge from both new and old tasks. $\mathbf{R}_{t+1}$ is obtained by weighted fusion of the projection of $\mathbf{Q}_{t}$ on $\mathbf{Q}_{t-1}$ with $\mathbf{R}_{t-1}$.
III-D Optimization
The loss function of our proposed method consists of two main parts: the cross-entropy loss and the dynamic distillation loss.
CLIP Loss: CLIP consists of an image encoder and a text encoder. The image x and the text z are separately input into the image encoder and text encoder, respectively, to obtain $\textbf{F}_{\text{image}}$ and $\textbf{F}_{\text{text}}$ respectively. In CLIP, z is obtained through manually crafted prompts, such as ”a good photo of [ ]”, where ”[ ]” is replaced by the class name of the test image. Therefore, the probability of predicting the test image x as class $y_{i}$ can be computed as:
$p\left(y_{i}\mid\textbf{x}\right)=\frac{e^{<\bm{F}_{\text{image}},\bm{F}_{\text{text}}^{y_{i}}>/\tau}}{\sum_{j=1}^{y_{t}}e^{<\bm{F}_{\text{image}},\bm{F}_{\text{text}}^{j}>/\tau}},$ (5)
[…]
After the Feature Calibration mechanism, we use the formula with $\bm{F}_{\text{fine-tuned}}$ to obtain the class probability vector and predict the class. The probability of predicting the test image x as class $y_{i}$ can be calculated as:
$p\left(y_{i}\mid\textbf{x}\right)=\frac{e^{<\alpha\textbf{F}_{\text{image }}+(1-\alpha)\Phi_{EFS}\left(\textbf{F}_{\text{image }}\right),\textbf{F}_{\text{text }}^{y_{i}}>/\tau}}{\sum_{j=1}^{y_{t}}e^{<\alpha\textbf{F}_{\text{image }}+(1-\alpha)\Phi_{EFS}(\textbf{F}_{\text{image }}),\textbf{F}_{\text{text }}^{j}>/\tau}},$ (6)
We employ the cross-entropy loss criterion to fine-tune the parameters of the adapter, formulated as follows:
$\mathcal{L}_{ce}(\textbf{y},\textbf{p})=-\sum_{i=1}^{n}y_{i}\log p_{i}.$ (7)
Dynamic Distill Loss: LwF [2] is the first successful case of applying knowledge distillation to CIL. Therefore, we directly incorporate this loss. The specific calculation is as follows:
$\mathcal{L}_{\text{distill }}=\sum_{j=1}^{\left|y_{t-1}\right|}-S_{k}\left(f^{t-1}(x)\right)\log S_{k}(f(x)).$ (8)
Due to the varying number of new and old classes in different incremental stages [26], we introduce a dynamic balancing parameter $\lambda$ and the overall loss is as follows: :
$\mathcal{L}=(1-\lambda)\mathcal{L}_{ce}+\lambda\mathcal{L}_{distill},$ (9)
where $\lambda=\left|\mathrm{y}_{\mathrm{t}-1}\right|/\left|\mathrm{y}_{\mathrm{t}}\right|$ represents the proportion of old classes among all classes. It increases as the incremental tasks evolve, indicating that the model focuses more on the old tasks.
TABLE I: Results on the ImageNet100 dataset.
Method ImageNet100 B0 ImageNet100 B50
Inc5 Inc10 Inc20 Inc5 Inc10
Avg Last Avg Last Avg Last Avg Last Avg Last
iCaRL [27] 69.00 50.90 74.10 58.50 78.10 65.20 57.30 44.40 60.70 44.70
LUCIR [28] 64.70 47.80 70.50 55.30 76.00 64.00 66.90 56.80 77.20 68.20
PODNet [29] 66.70 48.90 72.30 72.00 78.20 66.20 79.00 70.80 80.30 73.50
Cont-CLIP [14] 85.74 75.40 84.98 75.40 84.03 75.40 81.35 75.40 81.09 75.40
DualPrompt [15] 75.40 61.10 80.65 67.38 84.65 74.24 62.10 22.36 74.20 49.78
L2P++ [25] 75.43 62.10 80.51 67.22 84.12 73.70 62.00 22.15 74.11 49.46
ADAM [30] 85.78 75.72 85.84 76.40 85.85 77.08 84.90 78.58 84.60 78.58
SLCA [23] 78.40 63.36 78.63 59.92 84.08 71.08 86.47 72.22 86.26 71.18
CODA [16] 51.64 24.94 64.13 34.76 69.78 43.96 57.33 19.95 65.14 28.80
PROOF [31] 86.92 75.52 84.71 72.48 81.92 68.56 84.16 74.44 82.78 71.04
RARF [19] 87.59 79.87 87.51 80.23 86.72 80.10 86.53 80.16 86.36 80.22
Ours 88.04 78.32 88.19 80.32 87.62 81.00 85.50 79.70 85.86 80.40
IV Experiments
IV-A Experimental Setup
IV-A1 Datasets
The experiments were conducted on the CIFAR100 [32] and ImageNet100 [30] datasets. The CIFAR100 dataset contains 100 classes, with each class consisting of 600 color images of resolution 32 $\times$ 32 pixels. Each class contains 500 training samples and 100 test samples. ImageNet100 consists of samples from 100 classes, with each sample having a size of 224 $\times$ 224 pixels. Each class contains approximately 1300 training samples and 50 test samples. Our task division follows the two experimental configurations used in DER [32]: B0, where all classes are evenly distributed across different tasks, and B50, where the first task contains 50 classes (half of the dataset) and the remaining classes are evenly distributed across subsequent tasks. Inc denotes the number of incremental classes.
IV-A2 Implementation Details
We implemented the above method using PyTorch on an NVIDIA RTX 4090 GPU. We use the ViT-B/16 model as the backbone for CLIP. The optimizer is Adam, and each task is trained for 15 epochs. The initial learning rate is 0.001, and it is reduced by a factor of 0.1 at the 4th and 10th epochs. The batch sizes for the ImageNet100 and CIFAR100 datasets are 128 and 100, respectively. For each task, we generate approximately 2,000 samples by sampling from the Gaussian distribution of existing samples to simulate replay samples. The value of $\alpha$ is set to 0.8, and $\beta$ is set to 0.55 for ImageNet100 and 0.65 for CIFAR100. Under the B0 experimental configuration, we set up incremental settings with 5, 10, and 25 tasks. Under the B50 configuration, we set up incremental settings with 5 and 10 tasks.
IV-B Comparison with State-Of-The-Art Methods
Experimental results show that our method outperforms other existing methods in most cases across different class-incremental settings on the CIFAR100 and ImageNet100 datasets. The methods included in the comparison are mainly divided into two categories: traditional CIL methods and non-traditional CIL methods. Specifically, the results of the Continual-CLIP method refer to the zero-shot performance of the CLIP model. To ensure a fair comparison, all non-traditional CIL methods use the same CLIP pretrained weights.
TABLE II: Results on the CIFAR100 dataset.
Method CIFAR100 B0 CIFAR100 B50
Inc5 Inc10 Inc20 Inc5 Inc10
Avg Last Avg Last Avg Last Avg Last Avg Last
iCaRL [27] 61.20 44.90 65.30 50.90 71.10 59.60 58.60 49.50 65.10 56.00
LUCIR [28] 58.20 41.10 58.70 42.90 62.80 46.90 59.90 48.20 64.30 52.70
PODNet [29] 54.00 35.80 58.00 40.70 66.70 51.50 64.00 51.70 67.30 55.90
DualPrompt [15] 79.74 69.91 81.45 72.51 85.19 77.47 58.55 15.26 72.00 45.05
Cont-CLIP [14] 75.93 66.68 75.00 66.68 74.01 66.68 70.79 66.68 70.77 66.68
L2P++ [25] 79.18 68.67 81.90 73.08 84.39 77.37 58.57 18.04 76.51 48.52
ADAM [30] 70.18 58.12 80.53 65.50 77.28 67.89 83.38 76.94 83.21 76.94
CODA [16] 69.78 41.98 76.98 62.25 78.65 65.29 58.45 15.99 67.88 28.77
PROOF [31] 85.12 76.13 84.88 76.29 84.11 76.86 83.22 76.25 83.17 76.50
SLCA [23] 78.96 66.84 80.53 67.58 85.25 76.99 86.99 76.80 86.55 79.92
CIL-PVLM [18] 84.40 76.30 84.20 76.90 83.90 77.60 82.90 76.40 81.40 77.00
Ours 86.70 78.84 86.06 79.35 85.45 79.87 83.04 79.38 83.26 79.58
IV-B1 Results on ImageNet100
Table I shows the results of our method on ImageNet100 under five different incremental settings. It is clear that the non-traditional CIL method significantly outperforms traditional CIL methods, achieving a substantial performance gap. It indicates that CLIP holds great potential in the field of CIL, likely due to the positive impact of incorporating text features. Compared to the zero-shot performance of CLIP, our method achieves average incremental accuracy gains of 2.3%, 3.21%, 3.59%, 4.15%, and 4.77% on the B0Inc5, Inc10, Inc20, and B50Inc5, Inc10 incremental task settings, respectively. It suggests that our method is not solely reliant on the inherent generalization ability of CLIP but also the effectiveness of our proposed FC and PS techniques. Compared to DualPrompt [15] and CODA [16], our method performs better, demonstrating that optimizing and improving the text prompt space alone is insufficient. It is also necessary to enhance the use of visual features. Under the B0 experimental setting, our method achieved an average accuracy of 88.04%, 87.62%, and 88.19%, which are at least 0.45%, 0.68%, and 0.9% higher than those of other methods, showcasing the superiority of our approach. From the last metric, our method consistently ranks at the top, maintaining around 80% with slight fluctuations, whereas other method’s last values are much lower. It indicates that our method exhibits strong resistance to forgetting, confirming the rationality of incorporating parameter decomposition, fusion, and dynamic distillation ideas in our approach.
IV-B2 Results on CIFAR100
Table II shows the results of our method on CIFAR-100 under five different incremental settings. It is evident that the non-traditional CIL method outperforms traditional methods, with an even larger and more pronounced performance gap compared to its performance on the dataset ImageNet100. Compared to the zero-shot performance of CLIP, our method achieves average incremental accuracy gains of 10.77%, 11.06%, 11.44%, 12.25%, and 12.49% on the B0Inc5, Inc10, Inc20, and B50Inc5, Inc10 incremental task settings, respectively. These gains are nearly 2 to 3 times greater than those observed on the dataset ImageNet100, further validating the powerful capabilities of FC and PS. Under the B0 experimental setting, our method achieves an average accuracy of 86.70%, 86.06%, and 85.45% across different incremental task settings, which are at least 1.58%, 1.18%, and 0.26% higher than those of other methods. It demonstrates the excellence of our approach. The metric last of our method is consistently above 79.50% in various experimental settings, a level that is nearly unattainable for other methods. These highlight the significant advantage of our method in addressing the forgetting problem.
IV-C Ablation Studies
IV-C1 Module Ablation
TABLE III: Results of module ablation on the ImageNet100 B0 Inc10.
ID FC PS Distill Avg Last
1 84.98 75.40
2 85.82 73.78
3 88.16 80.02
4 88.19 80.32
Table III presents the results of ablation experiments on different modules of our proposed method. The FC module contributes a 0.84% improvement in accuracy, indicating that retaining the original CLIP features is crucial. It prevents the potential distortion of CLIP original features caused by direct fine-tuning. However, the last metric decreases by 1.62%. It suggests that FC somewhat exacerbates forgetting to a certain extent, while plays a significant role in learning new knowledge. The PS module, which uses QR decomposition to retain old knowledge, improves accuracy by 2.34% and the last metric by 6.24%. It demonstrates that the module not only positively impacts the learning of new knowledge but is also highly effective in addressing the forgetting problem. It indicates that decomposing and fusing the adapter’s parameter matrix is an efficient approach, balancing the learning of new knowledge while preserving old knowledge. The dynamic distillation module plays a more prominent role in mitigating the forgetting problem.
IV-C2 Effectiveness of Loss Function
In Eq. 9 , we primarily use cross-entropy loss and dynamic distillation loss, with the parameter $\lambda$ incorporated into the loss calculation. As shown in Table III, when we only use cross-entropy loss, the last accuracy of the method is 80.02%. When we introduce $\lambda$ and distillation loss, the last accuracy improves by 80.32%. Additionally, the $\lambda$ we introduce represents the ratio of old classes to all classes and is dynamically adjusted. When $\lambda$ is a fixed value, the accuracy decreases to some extent. It indicates that the dynamic design of $\lambda$, which considers the proportion of old and new classes, allows the loss function to adapt dynamically, leading to improved performance.
Figure 4: Results of performance sensitivity analysis on (a) $\alpha$ and (b) $\beta$ with ImageNet100 B0 Inc10 configuration.
IV-C3 Sensitivity Analysis on $\alpha$ and $\beta$
When preserving the original CLIP features, $\alpha$ represents the degree of preservation of the original features. It requires a trade-off between the CLIP features and the enhanced features. As shown in Fig. 4(a), the size of $\alpha$ has a certain impact on the accuracy, indicating that whether and to what extent the original CLIP features are preserved is crucial. Compared to $\alpha$, the size of $\beta$ has a more direct and profound impact on the accuracy of the model. An inappropriate $\beta$ can have a detrimental effect on the performance of model in Fig. 4(b), further emphasizing the importance of the decomposition and fusion of the adapter parameters.
V Conclusion
This paper proposes a novel CIL method called FCPS which integrates visual and textual features into the continual learning framework. A specific parameter update mechanism is designed to address the issue of feature distortion caused by directly fine-tuning vision-language models, effectively maintaining the stability and generalization of the model. Additionally, parameter update and integration approach is introduced for different task stages to mitigate catastrophic forgetting. Experimental results demonstrate the superiority of this method. Some parameter values were manually determined, and future work will focus on developing dynamic adaptive update methods while also exploring few-shot CIL to address the limited availability of new class data in real-world scenarios.