### v4

# Feature Calibration Enhanced Parameter Synthesis for CLIP-based Class-Incremental Learning

Abstract
Class-Incremental Learning (CIL) enables models to continuously learn new class knowledge while retaining previous classes, facilitating adaptation and evolution in dynamic, real-world environments. Traditional CIL methods primarily rely on visual features, which limits their effectiveness in complex, multimodal scenarios. In contrast, VLMs show promising potential for enhancing CIL by leveraging pre-trained knowledge and integrating multi-modal semantic cues such as text and vision. However, existing approaches struggle to mitigate catastrophic forgetting while preserving the generalization strengths of VLMs across diverse modalities. To address these challenges, we propose a Feature Calibration Enhanced Parameter Synthesis (FCPS) framework. Specifically, FCPS introduces a dynamic parameter adjustment mechanism that iteratively calibrates the contribution of original visual features to the final class decision, thus preserving the model’s intrinsic generalization capability across modalities. Simultaneously, parameter integration enables effective knowledge transfer, maintaining a balance between acquiring new class representations and preserving old knowledge. Experimental results on popular benchmarks (e.g., CIFAR100 and ImageNet100) validate the superiority of the proposed method.
Index Terms:
Class-Incremental Learning, Visual Language Learning, Knowledge Distillation, Feature Calibration.
I Introduction
Class-Incremental Learning (CIL) aims to incorporate the knowledge of new classes incrementally and build a universal classification model among all seen classes [1, 2, 3, 4, 5]. In other words, CIL should not only continuously acquire knowledge from the current task but also preserve the knowledge from former tasks, making it crucial for real-world applications [6, 7] that require adaptability. For example, autonomous driving systems must dynamically adapt to evolving road conditions, weather variations, traffic regulations, and driving patterns [8], while medical imaging systems face challenges from emerging diseases and evolving pathological presentations. These scenarios require CIL-enabled models that effectively integrate new information without revisiting previous data - a necessity for deployment on edge devices constrained by limited storage and computational capacity.
Figure 1: Illustration of existing CIL framework.
A primary challenge in CIL is to mitigate catastrophic forgetting [9], where the integration of new knowledge could overwrite previously acquired information in the process of continuously building a classifier for all classes. Effective solutions must balance stability—the retention of past knowledge—and plasticity—the adaptability to new tasks [10]. To address this, various approaches have been proposed, including replay-based[11], regularization-based [12], and parameter-isolated methods [13]. For instance, [11] proposes an efficient memory scheme based on the Resnet-18 to cope with the highly imbalanced data problem, while DNE [14] introduces a dynamic architecture that balances accuracy and model complexity based on Vision Transformer (ViT)[15] backbone. While these strategies have shown success with foundational vision backbones (as shown in Fig. 1(a)) like ResNet[16] and ViT[15], they focus solely on visual features, limiting their robustness in complex, multi-modal scenarios where integrating heterogeneous information sources is essential.
Recent advancements in incremental learning have explored Vision-Language Models (VLMs), such as ViT-BERT [17] and CLIP [18], which harness both visual and textual information to enhance performance. These pretrained models have learned rich feature representation from large-scale multi-modal data [19, 20], which can be used to bridge the knowledge gap between class information trained on different stages. By freezing CLIP’s pretrained backbone and leveraging its zero-shot capabilities, Continual-CLIP[21] demonstrates the pretained VLMs are well suited for CIL. Motivated by the zero-shot ability of CLIP, the follow-up methods propose different finetuning methods to mitigate catastrophic forgetting, consistently outperforming traditional vision-only methods [22].
Prompt-tuning[23, 24] and visual feature tuning [25] are two representative categories of CLIP-based approaches for incremental learning. Prompt-tuning employs task-specific or input-conditioned prompts to guide the model while preserving the original backbone, while visual feature tuning modifies the pretrained model to accommodate new tasks as illustrated in Fig. 1(c). Even though the previous methods can inject valuable pretrained knowledge, they still face challenges for CIL. First, prompt-tuning methods primarily focus on improving and enhancing the text prompt space in Fig. 1(b), overlooking the importance of effectively utilizing visual features. Second, visual feature tuning methods risk degrading its feature extraction ability based on its coverage of the original CLIP features. As shown in Fig. 2, the last-stage classification clusters of visual feature tuning method is not even seperated clearly compared to those of zero-shot CLIP. The recently proposed RAPF[26] is a visual feature tuning method that uses neighboring categories separation to enhance the representation of the feature space. However, the new and old knowledge is projected into one shared feature basis, limiting the ability to adapt to the shared feature space dynamically and adopting the CLIP features.
Figure 2: T-SNE visualization of (a) Continual-CLIP, (b) visual feature tuning and (c) Our method with CIFAR100 B0 Inc10 after the second task.
To address these issues, this paper proposes a novel CIL method called Feature Calibration Enhanced Parameter Synthesis (FCPS). We adopt the frozen encoder of CLIP as the foundational architecture for our approach. Considering the potential distortion of the pretrained knowledge of CLIP caused by fine-tuning, we designed Feature Calibration (FC) which uses the trainable layer to transform the CLIP features into new feature spaces. FC effectively utilizes and rationalizes the enhancement of the original visual features of CLIP while introducing specialized parameters to control the retention of the original visual features of CLIP precisely. It allows the model to maintain strong generalization capabilities while reducing the likelihood of overfitting. At the same time, we focus on whether the model can be dynamically adapted between old and new knowledge in the process of continuously building a classifier for all classes, rather than on the adaptation and improvement of textual cues. We introduce the Parameter Synthesis (PS) module, which effectively mitigates catastrophic forgetting by addressing the significant discrepancies in knowledge between adjacent tasks. PS integrates the adapter parameters of the current stage with those of the previous stage to initialize the adapter parameters for the next stage. With task evolution, the parameters of adjacent stages undergo decomposition and integration, alleviating the forgetting issue as tasks progress. Furthermore, we incorporate an additional dynamic distillation loss during the training process to constrain the changes of adapter parameters, further mitigating forgetting. The main contributions of this paper are as follows:
  • We propose a tuning-efficient VLM-based framework for CIL, which effectively solves the issue of the original feature of CLIP being distorted due to fine-tuning with a feature calibration module.
  • To tackle the catastrophic forgetting problem, we introduce the QR decomposition-based parameter synthesis, which captures parameter changes during task transitions and fuses the parameters constructed for different tasks.
  • Experimental results on CIFAR100 and ImageNet100 datasets show that our method achieves comparable performance to state-of-the-art methods.
The remainder of this paper is organized as follows. In Section II, we review related work on class-incremental learning and Vision-Language Models. Section III introduces the overall framework of our proposed method. We then elaborate on the design of the feature calibration and parameter synthesis modules, and clarify the optimization objective. In Section IV, we present comprehensive experimental results and compare FCPS with state-of-the-art methods. Finally, Section V concludes the paper by discussing limitations and potential future research directions.
II Related work
II-A Class-Incremental Learning
CIL aims to continuously incorporate new knowledge from new data while retaining, integrating and even optimizing old knowledge to prevent catastrophic forgetting [10] in the process of continuously building a classifier for all classes [27]. The current mainstream methods for CIL are divided into three main categories. The replay-based methods either save a portion of samples from previous tasks [28, 29, 30, 31] or generate synthetic samples with a generative model [32, 33, 34, 35, 36, 37, 38] in a cache. When training on new tasks, these samples are used alongside new samples to prevent forgetting previously learned knowledge. [29] selects samples with high predictive entropy and proximity to the decision boundary, aiming to enhance generalization by replaying these challenging examples. [28] estimates sample uncertainty through data augmentation, identifying instances with high prediction variance across multiple augmented versions. Regularization-based methods impose constraints on network parameter updates. Weight regularization focuses on preventing drift in weights associated with previous tasks [39, 12, 40] , while data regularization methods transfer knowledge from the old model to the new one through knowledge distillation [41, 42, 43] . Given that each parameter does not contribute equally to the task, parameter regularization methods seek to assess the importance of each parameter to the network and keep important parameters static to preserve previous knowledge. [40] examines the performance and stability of various network layers in CIL, revealing distinct characteristics. Shallow layers learn quickly but offer limited representational capacity, whereas deeper layers adapt more slowly yet provide stronger discriminative features. While knowledge distillation refers to building a mapping between a set of methods, which has been widely adopted in CIL in several ways. AFC [42] considers the importance of different feature maps for distillation. COIL [41] suggests the use of collaborative transport for bidirectional distillation, where both the semantic relationships between old and new models are exploited. Parameter isolation-based methods mitigate catastrophic forgetting by dynamically increasing the capacity of network. It enables the model to retain knowledge from previous tasks while expanding to learn new ones [44, 13, 45, 46, 47]. [46] proposes the Dynamically Expandable Representation (DER) method, which decouples feature adaptation from the classifier. For a new task, DER freezes the existing feature extractor, adds a new one, and concatenates features from all extractors for classification. [47] proposes a dynamic structure reorganization strategy that preserves old knowledge while learning new classes. When a new class arrives, side branches are added to the network blocks. The original weights are frozen, and the new branches are trained to learn the new class. Due to the excellent zero-shot capability of CLIP on single or multi-task [48, 49, 50, 51, 20], recent works [25, 23, 24] have shown the possibility of applying CLIP to incremental learning scenarios.
II-B Vision-Language Models
Large-scale Vision-Language Models such as CLIP [18], ALIGN [52], and Florence [53] have been pretrained on massive datasets containing image-text pairs, enabling them to learn a unified embedding space that bridges the gap between visual and textual modalities [54]. These models exhibit remarkable generalization and transfer capabilities, making them particularly effective for zero-shot learning across a wide range of downstream tasks [55] without the need for task-specific fine-tuning. In the context of incremental learning, a growing body of research has begun to explore how to adapt these powerful pretrained models to sequential tasks while minimizing catastrophic forgetting [56]. A recent trend involves introducing additional learnable components—such as prompts or lightweight modules—that allow the pretrained backbone to remain largely unchanged while adapting to new tasks. For instance, L2P [57] brings prompt-based learning into incremental learning by maintaining a shared pool of prompts that guide the model to adapt to each task dynamically. DualPrompt [22] proposes attaching complementary prompts to the pretrained backbone and formulates the objective as learning task-invariant and task-specific instructions. Most of these works focus on prompt adjustment, prefix adjustment and effective organization of text features but ignore the effective attention to visual features. This paper aims to fill this research gap by proposing a comprehensive solution that focuses on visual features to tune visual language models without suffering from forgetting.
III Methodology
III-A Overview
Traditional visual methods often struggle to make effective use of textual information [58], and fine-tuning methods based on CLIP have shown promising performance in multimodal settings. They are highly susceptible to catastrophic forgetting when applied to incremental learning scenarios. To address these limitations, we propose a novel method called FCPS, which effectively leverages the rich cross-modal knowledge encapsulated in CLIP while simultaneously addressing the challenge of forgetting. As shown in Fig.3, FCPS enhances performance through three core components that collectively aim to balance knowledge retention and task comprehensiveness. The first component FC is introduced to preserve and optimize the pretrained CLIP representations. To reduce the risk of distorting the pretrained CLIP knowledge caused by direct fine-tuning, FC is designed to refine the extracted visual features in a lightweight and stable manner. This module enhances the visual embeddings without altering the core CLIP encoder, ensuring that the foundational multimodal knowledge is preserved. To further maintain the generalization capability of the model while preventing overfitting to the current task, we introduce a parameter $\alpha$ after the calibration stage. This parameter determines the extent to which the original image embeddings from the CLIP Image Encoder are retained. By maintaining a proportion of these original embeddings, the model can better preserve the semantics of the pretrained representation space, leading to improved robustness across diverse tasks. The second core component, PS, is designed to address the issue of forgetting. It effectively integrates the current adapter parameters with those from the previous stage, serving as the initial parameters for the next stage’s adapter. This progressive update mechanism allows knowledge from earlier tasks to be preserved and integrated into future learning. As tasks evolve, PS accumulates experience by continuously adapting and refining its parameters, ensuring that knowledge from old classes is retained even as new ones are introduced. This design effectively mitigates catastrophic forgetting in the process of continuously building a classifier for all classes. Finally, FCPS introduces Dynamic Distillation, a module responsible for transferring knowledge between the current and previous tasks. This component acts as a bridge between stages, enhancing continuity and semantic consistency throughout the learning process. In summary, FCPS offers a comprehensive and systematic approach to incremental learning with VLMs. By preserving cross-modal knowledge, promoting effective parameter reuse, and introducing a distillation mechanism, FCPS achieves a strong balance between performance on new tasks and retention of previously acquired knowledge.
Figure 3: The framework of the proposed feature calibration enhanced parameter synthesis.
III-B Feature Calibration
Given the powerful zero-shot capabilities of CLIP and its remarkable transferability across a wide range of downstream tasks, we introduce a feature calibration mechanism to better align task-specific requirements with the pretrained CLIP representations. This mechanism is carefully designed to integrate the discriminative features required for a specific task with the generalizable and robust features extracted by CLIP, thereby enhancing the model’s adaptability without sacrificing its generalization performance. As depicted in Fig. 3, the image features $\mathbf{F}_{\text{fine-tuned}}$ are obtained by first passing input images through the frozen CLIP image encoder, followed by a lightweight adapter module. This adapter enables subtle yet effective transformation of the visual features, adapting them to the current task while minimizing the risk of disrupting the original pretrained knowledge. In parallel, label features $\mathbf{F}_{\text{text}}$ are generated by processing the textual class labels through a prompt generator, and then passing them through the CLIP text encoder. The generated features capture semantic class-level information, aligned in the same embedding space as the image features. Finally, classification is performed by computing the similarity between the calibrated image features and the class text features.
To further improve the representational capacity of the visual features while preserving compatibility with the pretrained CLIP framework, we propose a trainable adaptive module named Enhanced Feature Synthesis (EFS). This module is specifically designed to enhance feature quality through a deeper transformation process, while maintaining the original dimensionality of the input features to ensure seamless integration with downstream components. The EFS consists of four sequential linear transformation layers that enable the network to capture more complex relationships within the feature space. In addition to these transformations, the module incorporates attention mechanism techniques to dynamically recalibrate the feature responses. Importantly, the feature dimensionality remains unchanged before and after passing through the EFS. This design ensures that the enhanced features produced by EFS remain fully compatible with the CLIP feature space and can be directly used for similarity computation. To retain the original characteristics of CLIP, feature calibration is conducted as follows:
$\textbf{F}_{\text{fine-tuned}}=(1-\alpha)\Phi_{EFS}\left(\textbf{F}_{\text{image }}\right)+\alpha\textbf{F}_{\text{image}},$ (1)
where $\alpha$ is a parameter that balances the enhanced features with the original CLIP knowledge. The first term represents the adaptive learning of knowledge from new categories in the new task, while the second term is specifically designed to retain the original knowledge from CLIP.
III-C Parameter Synthesis
To address the challenge of catastrophic forgetting in incremental learning, we introduce parameter synthesis—a mechanism that dynamically integrates adapter parameters from the current learning stage with those preserved from previous stages. This integration enables the progressive fusion of new and old knowledge, ensuring that the model can incrementally construct a unified classifier capable of handling all previously encountered classes. By retaining and reusing historical parameters, PS promotes long-term memory retention while simultaneously adapting to novel information, thereby mitigating the risk of performance degradation due to forgetting.
In many traditional weight update schemes, newly acquired information often overwrites past knowledge, leading to a well-known issue of catastrophic forgetting. This problem is particularly pronounced in incremental learning scenarios, where the model is required to adapt continuously to a sequence of tasks without access to the full dataset from earlier stages. To further combat this issue, we propose a strategy based on QR decomposition. Specifically, during each incremental update, the model’s weight matrix is decomposed into an orthogonal matrix $\mathbf{Q}$ and an upper triangular matrix $\mathbf{R}$, allowing us to restructure and recombine the parameter space in a manner that preserves essential features of previously learned tasks. The use of QR decomposition offers a mathematically grounded approach to separating and retaining knowledge components from earlier stages, while facilitating the smooth integration of new task-specific information. By recombining these components in a controlled way, the model can maintain a balanced representation across tasks, effectively reducing interference and preserving critical knowledge.
Specifically, QR decomposition is employed to decompose a target weight matrix into two constituent matrices: an orthogonal matrix $\mathbf{Q}$ and an upper triangular matrix $\mathbf{R}$, such that the original matrix is represented as $\mathbf{W}=\mathbf{Q}\mathbf{R}$. In this decomposition, the matrix $\mathbf{Q}$ contains orthonormal basis vectors that span the column space of $\mathbf{W}$, while $\mathbf{R}$ captures the scaling and linear combinations required to reconstruct the original representation. As shown in Fig.4, We apply QR decomposition to the old and new weight matrices $\{\mathbf{W}_{t-1},\mathbf{W}_{t}\}$:
$\mathbf{W}_{t-1}=\mathbf{Q}_{t-1}\mathbf{R}_{t-1},\mathbf{W}_{t}=\mathbf{Q}_{t}\mathbf{R}_{t}.$ (2)
To achieve a controllable and effective integration of knowledge across tasks, we further introduce a weight coefficient $\beta$ to balance the contributions of the current and previous stages. This coefficient regulates the extent to which the newly acquired task information should influence the model, relative to the preserved knowledge from earlier tasks. Specifically, after performing QR decomposition on both the current and previous adapter weight matrices, we obtain corresponding $\mathbf{Q}$ and $\mathbf{R}$ components. We then linearly combine these components using the balancing coefficient $\beta$. The final fused weight matrix, which becomes the initial weight matrix for the model’s adapter in the next stage, is updated as:
$\mathbf{W}=\mathbf{Q}_{t-1}\mathbf{R}_{t+1},$ (3)
$\mathbf{R}_{t+1}=\left[\mathbf{Q}_{t-1}^{T}\otimes\mathbf{Q}_{t}\cdot\beta\right]\oplus\left[\mathbf{R}_{t-1}\cdot(1-\beta)\right],$ (4)
where $\beta$ is adjusted to maximize the retention of knowledge from both new and old tasks. $\mathbf{R}_{t+1}$ is obtained by weighted fusion of the projection of $\mathbf{Q}_{t}$ on $\mathbf{Q}_{t-1}$ with $\mathbf{R}_{t-1}$.
Through QR decomposition, we can not only avoid the parameter drift caused by direct fine-tuning but also maintain the stability of the parameters in each task iteration, thus alleviating the problem of catastrophic forgetting. The advantage of this method is that QR decomposition preserves the structural information of the old weights and effectively fuses the learning of new tasks with the knowledge from old tasks, maintaining the overall generalization ability and knowledge transfer capability of the model.
Figure 4: The parameter synthesis process based on QR decomposition.
III-D Optimization
The loss function of our proposed method consists of two main parts: the cross-entropy loss and the dynamic distillation loss in Fig. 5.
CLIP Loss: CLIP consists of an image encoder and a text encoder. The image x and the text z are separately input into the image encoder and text encoder, respectively, to obtain $\textbf{F}_{\text{image}}$ and $\textbf{F}_{\text{text}}$ respectively. In CLIP, z is obtained through manually crafted prompts, such as ”a good photo of [ ]”, where ”[ ]” is replaced by the class name of the test image. Therefore, the probability of predicting the test image x as class $y_{i}$ can be computed as:
$p\left(y_{i}\mid\textbf{x}\right)=\frac{e^{<\bm{F}_{\text{image}},\bm{F}_{\text{text}}^{y_{i}}>/\tau}}{\sum_{j=1}^{y_{t}}e^{<\bm{F}_{\text{image}},\bm{F}_{\text{text}}^{j}>/\tau}},$ (5)
[…]
After the FC mechanism, we use the formula with $\bm{F}_{\text{fine-tuned}}$ to obtain the class probability vector and predict the class. The probability of predicting the test image x as class $y_{i}$ can be calculated as:
$p\left(y_{i}\mid\textbf{x}\right)=\frac{e^{<\alpha\textbf{F}_{\text{image }}+(1-\alpha)\Phi_{EFS}\left(\textbf{F}_{\text{image }}\right),\textbf{F}_{\text{text }}^{y_{i}}>/\tau}}{\sum_{j=1}^{y_{t}}e^{<\alpha\textbf{F}_{\text{image }}+(1-\alpha)\Phi_{EFS}(\textbf{F}_{\text{image }}),\textbf{F}_{\text{text }}^{j}>/\tau}},$ (6)
We employ the cross-entropy loss criterion to fine-tune the parameters of the adapter, formulated as follows:
$\mathcal{L}_{ce}(\textbf{y},\textbf{p})=-\sum_{i=1}^{n}y_{i}\log p_{i}.$ (7)
Dynamic Distill Loss: LwF [9] is the first successful case of applying knowledge distillation to CIL. Therefore, we directly incorporate this loss. The specific calculation is as follows:
$\mathcal{L}_{\text{distill }}=\sum_{j=1}^{\left|y_{t-1}\right|}-S_{k}\left(f^{t-1}(x)\right)\log S_{k}(f(x)).$ (8)
Due to the varying number of new and old classes in different incremental stages [59], we introduce a dynamic balancing parameter $\lambda$ and the overall loss is as follows: :
$\mathcal{L}=(1-\lambda)\mathcal{L}_{ce}+\lambda\mathcal{L}_{distill},$ (9)
where $\lambda=\left|\mathrm{y}_{\mathrm{t}-1}\right|/\left|\mathrm{y}_{\mathrm{t}}\right|$ represents the proportion of old classes among all classes. It increases as the incremental tasks evolve, indicating that the model focuses more on the old tasks.
Figure 5: The loss function construction for our proposed method.
TABLE I: Results on the ImageNet100 dataset.
Method ImageNet100 B0 ImageNet100 B50
Inc5 Inc10 Inc20 Inc5 Inc10
Avg Last Avg Last Avg Last Avg Last Avg Last
iCaRL [60] 69.00 50.90 74.10 58.50 78.10 65.20 57.30 44.40 60.70 44.70
LUCIR [61] 64.70 47.80 70.50 55.30 76.00 64.00 66.90 56.80 77.20 68.20
PODNet [62] 66.70 48.90 72.30 72.00 78.20 66.20 79.00 70.80 80.30 73.50
Cont-CLIP [21] 85.74 75.40 84.98 75.40 84.03 75.40 81.35 75.40 81.09 75.40
DualPrompt [22] 75.40 61.10 80.65 67.38 84.65 74.24 62.10 22.36 74.20 49.78
L2P++ [57] 75.43 62.10 80.51 67.22 84.12 73.70 62.00 22.15 74.11 49.46
ADAM [63] 85.78 75.72 85.84 76.40 85.85 77.08 84.90 78.58 84.60 78.58
SLCA [50] 78.40 63.36 78.63 59.92 84.08 71.08 86.47 72.22 86.26 71.18
CODA [23] 51.64 24.94 64.13 34.76 69.78 43.96 57.33 19.95 65.14 28.80
PROOF [64] 86.92 75.52 84.71 72.48 81.92 68.56 84.16 74.44 82.78 71.04
RARF [26] 87.59 79.87 87.51 80.23 86.72 80.10 86.53 80.16 86.36 80.22
Ours 88.04 78.32 88.19 80.32 87.62 81.00 85.50 79.70 85.86 80.40
IV Experiments
IV-A Experimental Setup
IV-A1 Datasets
The experiments were conducted on the CIFAR100 [46] and ImageNet100 [63] datasets. The CIFAR100 dataset contains 100 classes, with each class consisting of 600 color images of resolution 32 $\times$ 32 pixels. Each class contains 500 training samples and 100 test samples. ImageNet100 consists of samples from 100 classes, with each sample having a size of 224 $\times$ 224 pixels. Each class contains approximately 1300 training samples and 50 test samples. Our task division follows the two experimental configurations used in DER [46]: B0, where all classes are evenly distributed across different tasks, and B50, where the first task contains 50 classes (half of the dataset) and the remaining classes are evenly distributed across subsequent tasks. Inc denotes the number of incremental classes.
IV-A2 Evaluation Metrics
In the t-th incremental phase, the incremental accuracy $A_{t}$ denotes the classification accuracy of the current model over all encountered classes up to that point. To facilitate fair comparison across different methods, the average incremental accuracy is computed as the mean accuracy across all incremental phases: $Avg=(1/T)\sum_{i=1}^{T}A_{i}$, where $T$ is the total number of incremental learning phases. The $Last$ metric represents the overall accuracy after completing the final task.
IV-A3 Implementation Details
We implemented the above method using PyTorch on an NVIDIA RTX 4090 GPU. We use the ViT-B/16 model as the backbone for CLIP. The optimizer is Adam, and each task is trained for 15 epochs. The initial learning rate is 0.001, and it is reduced by a factor of 0.1 at the 4th and 10th epochs. The batch sizes for the ImageNet100 and CIFAR100 datasets are 128 and 100, respectively. For each task, we generate approximately 2,000 samples by sampling from the Gaussian distribution of existing samples to simulate replay samples. The value of $\alpha$ is set to 0.8, and $\beta$ is set to 0.55 for ImageNet100 and 0.65 for CIFAR100. Under the B0 experimental configuration, we set up incremental settings with 5, 10, and 25 tasks. Under the B50 configuration, we set up incremental settings with 5 and 10 tasks. The experimental results of the comparison methods are obtained by running the respective public codes.
IV-B Comparison with State-Of-The-Art Methods
Experimental results show that our method outperforms other existing methods in most cases across different class-incremental settings on the CIFAR100 and ImageNet100 datasets. The methods included in the comparison are mainly divided into two categories: traditional CIL methods and non-traditional CIL methods. Specifically, the results of the Continual-CLIP method refer to the zero-shot performance of the CLIP model. To ensure a fair comparison, all non-traditional CIL methods use the same CLIP pretrained weights.
TABLE II: Results on the CIFAR100 dataset.
Method CIFAR100 B0 CIFAR100 B50
Inc5 Inc10 Inc20 Inc5 Inc10
Avg Last Avg Last Avg Last Avg Last Avg Last
iCaRL [60] 61.20 44.90 65.30 50.90 71.10 59.60 58.60 49.50 65.10 56.00
LUCIR [61] 58.20 41.10 58.70 42.90 62.80 46.90 59.90 48.20 64.30 52.70
PODNet [62] 54.00 35.80 58.00 40.70 66.70 51.50 64.00 51.70 67.30 55.90
DualPrompt [22] 79.74 69.91 81.45 72.51 85.19 77.47 58.55 15.26 72.00 45.05
Cont-CLIP [21] 75.93 66.68 75.00 66.68 74.01 66.68 70.79 66.68 70.77 66.68
L2P++ [57] 79.18 68.67 81.90 73.08 84.39 77.37 58.57 18.04 76.51 48.52
ADAM [63] 70.18 58.12 80.53 65.50 77.28 67.89 83.38 76.94 83.21 76.94
CODA [23] 69.78 41.98 76.98 62.25 78.65 65.29 58.45 15.99 67.88 28.77
PROOF [64] 85.12 76.13 84.88 76.29 84.11 76.86 83.22 76.25 83.17 76.50
CIL-PVLM [25] 84.40 76.30 84.20 76.90 83.90 77.60 82.90 76.40 81.40 77.00
RARF [26] 86.61 78.68 86.01 79.27 85.65 79.82 83.19 79.45 83.06 79.85
Ours 86.70 78.84 86.06 79.35 85.45 79.87 83.04 79.38 83.26 79.58
IV-B1 Results on ImageNet100
Table I presents the performance of our proposed method on the ImageNet100 dataset under five distinct incremental learning settings. It is evident from the results that our approach, which deviates from conventional CIL paradigms, significantly outperforms traditional methods. The observed performance gap is substantial and consistent across all settings, highlighting the effectiveness of integrating CLIP into the CIL framework. This performance boost is largely attributed to the incorporation of text features, which enhance the semantic alignment between visual and textual modalities. These results underscore the promising potential of CLIP in incremental learning scenarios, where maintaining performance across sequential tasks is notoriously challenging. Compared to the zero-shot performance of CLIP, our method achieves average incremental accuracy gains of 2.3%, 3.21%, 3.59%, 4.15%, and 4.77% on the B0Inc5, Inc10, Inc20, and B50Inc5, Inc10 incremental task settings, respectively. It suggests that our framework does not rely solely on the strong generalization capabilities of pre-trained CLIP representations. Instead, the gains reflect the effectiveness of our novel FC and PS strategies, which are specifically designed to address the challenges of catastrophic forgetting and representation shift in CIL. Compared to DualPrompt [22] and CODA [23], our method consistently achieves superior results. While DualPrompt and CODA focus primarily on optimizing the textual prompt space, our findings suggest that improvements in prompt engineering alone are insufficient for robust CIL performance. Our approach leverages both textual and visual modalities by enhancing visual feature utilization, demonstrating that a more holistic integration of multimodal representations is critical. Under the B0 experimental setting, our method achieve average accuracy of 88.04%, 87.62%, exceeding the next-best methods by margins of at least 0.45%, 0.68%, and 0.9%, respectively. These consistent gains further validate the robustness and superiority of our proposed techniques. Additionally, when evaluating the performance based on the final task accuracy—a critical indicator of a model’s resistance to forgetting—our method maintains a performance level around 80%. In contrast, other methods exhibit significantly lower final accuracies, indicating greater susceptibility to catastrophic forgetting. These findings confirm that our method offers a well-balanced and effective solution for CIL. The integration of parameter decomposition, feature fusion, and dynamic distillation not only mitigates the drawbacks of previous approaches but also introduces a resilient framework that adapts well to evolving class distributions.
IV-B2 Results on CIFAR100
Table II presents the performance evaluation of our proposed method on the CIFAR-100 dataset under five distinct incremental learning scenarios. As shown, non-traditional CIL approach significantly outperforms existing traditional methods, exhibiting an even more pronounced performance gap compared to the results observed on the ImageNet100 dataset. This substantial improvement underscores the generalizability and robustness of our approach across different datasets, including those with smaller resolution and less semantic complexity like CIFAR-100. Compared to the zero-shot performance of CLIP, our method achieves average incremental accuracy gains of 10.77%, 11.06%, 11.44%, 12.25%, and 12.49% on the B0Inc5, Inc10, Inc20, and B50Inc5, Inc10 incremental task settings, respectively. These gains are approximately two to three times greater than those observed on ImageNet100, further confirming the effectiveness of the proposed FC and PS mechanisms. Under the B0 experimental setting, our method achieve average accuracy of 86.70%, 86.06%, and 85.45% across different incremental task settings. These results are consistently higher than those achieved by competing methods, providing strong evidence for the robustness and superiority of our approach. Furthermore, the last metric, which indicates the final task performance and serves as a key measure of a model’s resilience to catastrophic forgetting, remains consistently above 79.30% across all experimental settings. In contrast, other methods are nearly unattainable to reach this level, often experiencing significant performance degradation over time. This consistent and high final performance highlights the effectiveness of our method in mitigating the forgetting problem—a central challenge in CIL. Overall, the outstanding performance of our method on CIFAR-100, demonstrates that our approach not only scales well across datasets but also provides a powerful and general framework for CIL.
IV-C Ablation Studies
IV-C1 Module Ablation
TABLE III: Results of module ablation on the ImageNet100 B0 Inc10.
FC PS Distill $A_{1}$ $A_{2}$ $A_{3}$ $A_{4}$ $A_{5}$ $A_{6}$ $A_{7}$ $A_{8}$ $A_{9}$ $A_{10}$ (Last) Avg
1 97.20 95.20 91.93 88.75 83.56 81.80 81.00 78.89 76.10 75.40 84.98
2 96.6 97.00 92.43 90.80 83.64 84.13 82.27 80.97 76.56 73.78 85.82
3 96.80 96.60 93.87 91.70 87.20 85.52 85.03 84.60 80.27 80.02 88.16
4 96.80 96.10 94.27 91.50 87.16 85.43 85.09 84.82 80.42 80.32 88.19
Table III presents the results of our ablation study, which evaluates the individual contributions of each module in our proposed method. Overall, our method demonstrates stable performance throughout all incremental stages. Meanwhile, the table shows the role of FC, PS, and distillation in enhancing model performance and mitigating catastrophic forgetting. The inclusion of the FC module leads to a 0.84% increase in average incremental accuracy, demonstrating that preserving the integrity of the original CLIP features is highly beneficial. FC prevents the distortion of pre-trained visual embeddings that often results from direct fine-tuning, thereby retaining the strong generalization capacity of CLIP. However, we observe a 1.62% drop in the last metric when FC is used alone, indicating that while it facilitates the acquisition of new knowledge, it may introduce a trade-off in terms of long-term memory retention. This suggests that FC enhances short-term task performance but slightly exacerbates forgetting over time if used in isolation. PS shows a more substantial contribution. It uses QR decomposition to retain old knowledge, improves accuracy by 2.34% and the last metric by 6.24%. These results clearly indicate that PS not only supports effective learning of new tasks but also plays a critical role in preserving previously learned knowledge. This module leverages the orthogonality principle in QR decomposition to facilitate better knowledge retention across tasks. The improvement in the last metric suggests that decomposing and fusing the adapter’s parameter matrix is a highly effective strategy, achieving a delicate balance between plasticity (learning new tasks) and stability (retaining old knowledge). The dynamic distillation module plays a more prominent role in mitigating the forgetting problem. By adaptively distilling knowledge from earlier task models and dynamically updating the student model, this module reduces the drift of representations and promotes consistency across learning stages. Its contribution becomes especially evident in the sustained high values of the last metric when combined with other modules, underscoring its vital role in preventing catastrophic forgetting. Overall, the ablation study validates the effectiveness of each individual module in our framework. While FC aids in preserving feature quality and learning new information, PS ensures that old knowledge is not overwritten, and dynamic distillation offers a strong regularization mechanism for incremental learning. The synergy among these components forms the foundation of our robust and adaptive CIL framework.
IV-C2 Effectiveness of Loss Function
In Eq.9, our loss function is composed of two key components: the standard cross-entropy loss, which facilitates the learning of new class information, and a dynamic distillation loss, which serves to preserve knowledge from previously learned tasks. These two components are weighted by a scalar parameter $\lambda$, which plays a crucial role in balancing the influence of old and new class information during training. As evidenced by the ablation results in TableIII, this design significantly impacts the model’s ability to mitigate forgetting and retain prior knowledge. When only the cross-entropy loss is used—without incorporating any distillation mechanism—the final task accuracy (last) is 80.02%. This baseline reflects the typical performance of a model focused solely on learning new information without explicit mechanisms to retain prior knowledge. However, when we introduce the distillation loss alongside the cross-entropy loss, and dynamically modulate their relative contributions using the parameter $\lambda$, the last metric improves to 80.32%. Although the numerical gain may appear modest, it demonstrates that even a small degree of incremental knowledge transfer can contribute meaningfully to long-term retention in CIL. The parameter $\lambda$ is not static, rather, it is dynamically computed based on the ratio of old classes to the total number of classes seen so far in the incremental learning process. This adaptive strategy ensures that as the task progresses and the balance between old and new classes changes, the loss function remains sensitive to the shifting importance of knowledge retention versus acquisition. In contrast, when $\lambda$ is fixed across all learning stages, the model’s accuracy decreases. This drop illustrates the limitation of static weighting schemes in dynamic learning environments. The results clearly indicate that the dynamic design of $\lambda$ contributes to better performance by providing an adaptive mechanism that reflects the class distribution at each stage. As the number of classes increases, and the potential for forgetting previously learned classes grows, the dynamic $\lambda$ ensures that the distillation loss is appropriately weighted to preserve old knowledge. Conversely, in early stages with fewer old classes, the model is allowed greater flexibility to focus on new knowledge acquisition. This adaptability allows our loss formulation to strike an effective balance between stability and plasticity, which is essential in incremental learning settings. Overall, the incorporation of a dynamically adjusted $\lambda$ improves the effectiveness of the distillation process. This design choice further reinforces the robustness and adaptability of our proposed framework in addressing the challenges of catastrophic forgetting.
Figure 6: Results of performance sensitivity analysis on (a) $\alpha$ and (b) $\beta$ with ImageNet100 B0 Inc10 configuration.
IV-C3 Sensitivity Analysis on $\alpha$ and $\beta$
In our framework, preserving the original CLIP features is a key design consideration, and we introduce the parameter $\alpha$ to control the degree of feature preservation. Specifically, $\alpha$ serves as a balancing coefficient that governs the trade-off between the original frozen CLIP features and the newly enhanced features derived from task-specific adaptations. As illustrated in Fig. 6(a), variations in the value of $\alpha$ yield noticeable changes in model accuracy, highlighting the sensitivity of the model’s performance to this balance. This observation underscores the importance of carefully calibrating how much of the pre-trained knowledge is retained versus how much is adapted to new tasks. When $\alpha$ is too small, the model may underutilize the rich semantic representations encoded in CLIP, leading to suboptimal generalization. Conversely, when $\alpha$ is too large, the model may become overly reliant on static features, limiting its ability to adapt to new class distributions in the incremental learning setting. Thus, selecting an appropriate value for $\alpha$ is essential to achieving a balanced integration of pre-trained and task-specific knowledge. In contrast to $\alpha$, the parameter $\beta$, which controls the influence of the decomposed and fused adapter parameters, has an even more direct and pronounced effect on model performance. As depicted in Fig. 6(b), improper tuning of $\beta$ can significantly degrade accuracy, revealing the sensitivity of the model to the fusion mechanism applied to the adapter’s parameters. This further emphasizes the critical role of parameter decomposition and fusion in our approach. The adapter module, which operates as a lightweight, trainable bridge between frozen CLIP representations and the downstream incremental tasks, relies heavily on the proper scaling and integration of its components. $\beta$ modulates the contribution of these components, and an ill-suited value can disrupt the representational harmony between old and new knowledge, ultimately hindering the model’s capacity to learn incrementally without forgetting. The experiments shown in Fig. 6 collectively reveal that both $\alpha$ and $\beta$ are vital hyperparameters, but they influence different aspects of the learning process. While $\alpha$ governs the trade-off between stability (retaining CLIP’s pre-trained knowledge) and plasticity (learning task-specific enhancements), $\beta$ directly affects the internal representational learning within the adapter. Optimal performance is achieved only when both parameters are carefully tuned, enabling a synergistic interaction between feature preservation and adaptable parameterization.
V Conclusion
This paper proposes a novel CIL framework called FCPS, which integrates visual and textual features within a incremental learning paradigm. To address the feature distortion caused by directly fine-tuning VLMs such as CLIP, FCPS introduces FC to preserve the original semantic structure of pre-trained features. This design helps maintain the model’s generalization ability and stability across tasks. In addition, FCPS employs a stage-specific parameter integration strategy to mitigate catastrophic forgetting. By adaptively decomposing and fusing parameters at different task stages, the model can effectively balance the learning of new knowledge with the retention of previous knowledge. This approach enables FCPS to achieve strong performance in both incremental accuracy and long-term knowledge retention. Experimental results on benchmarks like ImageNet100 and CIFAR-100 demonstrate that FCPS outperforms traditional and recent CIL methods, confirming the advantages of multi-modal feature integration and the proposed training strategies. While some hyperparameters in FCPS were manually set, future work will focus on developing adaptive parameter tuning methods. Additionally, we aim to extend FCPS to few-shot CIL scenarios, which better reflect real-world conditions where new class data is limited.